# Diffusion State-Guided Projected Gradient for Inverse Problems

### v1

@@ 4 Results
4.1 Experimental Setup
We evaluate the performance of DiffStateGrad applied to three SOTA diffusion methods of PSLD (Rout et al., 2023), ReSample (Song et al., 2023a), and DAPS (Zhang et al., 2024). These methods span both latent solvers (PSLD and ReSample) and pixel-based solvers (DAPS). We evaluate their performance based on key quantitative metrics, including LPIPS (Learned Perceptual Image Patch Similarity), PSNR (Peak Signal-to-Noise Ratio), and SSIM (Structural Similarity Index) (Wang et al., 2004). We demonstrate the effectiveness of DiffStateGrad on two datasets: a) the FFHQ $256\times 256$ validation dataset (Karras et al., 2021), and b) the ImageNet $256\times 256$ validation dataset (Deng et al., 2009). For pixel-based experiments, we use (i) the pre-trained diffusion model from (Chung et al., 2023) for the FFHQ dataset, and (ii) the pre-trained model from (Dhariwal & Nichol, 2021) for the ImageNet dataset. For latent diffusion experiments, we use (i) the unconditional LDM-VQ-4 model trained on FFHQ (Rombach et al., 2022) for the FFHQ dataset, and (ii) the Stable Diffusion v1.5 (Rombach et al., 2022) model for the ImageNet dataset.
We consider both linear and nonlinear inverse problems for natural images. For evaluation, we sample a fixed set of $100$ images from the FFHQ and ImageNet validation sets. Images are normalized to the range $[0,1]$. We use the default settings for PSLD, ReSample, and DAPS for all experiments (see Appendix C for more details).
For linear inverse problems, we consider (1) box inpainting, (2) random inpainting, (3) Gaussian deblur, (4) motion deblur, and (5) super-resolution. In the box inpainting task, a random $128\times 128$ box is used, while the random inpainting task employs a $70\%$ random mask. Gaussian and motion deblurring tasks utilize kernels of size $61\times 61$, with standard deviations of $3.0$ and $0.5$, respectively. For super-resolution, images are downscaled by a factor of $4$ using a bicubic resizer. For nonlinear inverse problems, we consider (1) phase retrieval, (2) nonlinear deblur, and (3) high dynamic range (HDR). For phase retrieval, we use an oversampling rate of 2.0, and due to the instability and non-uniqueness of reconstruction, we adopt the strategy from DPS (Chung et al., 2023) and DAPS (Zhang et al., 2024), generating four separate reconstructions and reporting the best result. Like DAPS (Zhang et al., 2024), we normalize the data to lie in the range [0, 1] before applying the discrete Fourier transform. For nonlinear deblur, we use the default setting from (Tran et al., 2021). For HDR, we use a scale factor of 2. We note that PSLD is not designed to handle nonlinear inverse problems.
@@ 4.1 Experimental Setup
(b) Gaussian Deblur.
Figure 7: Robustness of DiffStateGrad to measurement noise. We evaluate the performance of DiffStateGrad-PSLD and PSLD-DAPS with their respective counterparts on different tasks across a range of measurement noise levels using FFHQ $256\times 256$.
4.2 Robustness
Figure 6 exhibits the sensitivity of PSLD to the choice of MG step size; the performance of PSLD significantly deteriorates when a relatively large MG step size is used, leading to poor results across all tasks. In contrast, DiffStateGrad-PSLD shows superior robustness, maintaining high performance over a wide range of MG step sizes. Figure 7 demonstrates the robustness of DiffStateGrad methods compared to their non-DiffStateGrad counterparts when faced with increasing measurement noise. For both inpainting and Gaussian deblur tasks, the performance of DAPS and PSLD deteriorates significantly as noise levels rise. In contrast, DiffStateGrad-DAPS and DiffStateGrad-PSLD exhibit superior resilience across the range of noise levels tested.
Table 2: Performance comparison for inpainting tasks on FFHQ 256 $\times$ 256.
Method Inpaint (Box) Inpaint (Random)
@@ 4.2 Robustness
DAPS 0.136 (0.027) 0.806 (0.028) 24.57 (2.61) 0.130 (0.022) 0.829 (0.022) 30.79 (1.60)
DiffStateGrad-DAPS (ours) 0.113 (0.020) 0.849 (0.029) 24.78 (2.62) 0.099 (0.020) 0.887 (0.023) 32.04 (1.79)
@@
PSLD 0.158 (0.023) 0.819 (0.031) 24.22 (2.80) 0.246 (0.053) 0.809 (0.049) 29.05 (1.98)
DiffStateGrad-PSLD (ours) 0.092 (0.019) 0.880 (0.028) 24.32 (2.78) 0.165 (0.035) 0.898 (0.024) 31.68 (1.74)
ReSample 0.198 (0.032) 0.807 (0.036) 19.91 (1.78) 0.115 (0.028) 0.892 (0.030) 31.27 (2.35)
DiffStateGrad-ReSample (ours) 0.156 (0.027) 0.841 (0.032) 23.59 (2.38) 0.106 (0.023) 0.913 (0.023) 31.91 (1.84)
Table 3: Performance comparison for deblurring and super-resolution tasks on FFHQ 256 $\times$ 256.
Method Gaussian deblur Motion deblur SR (x4)
[…]
Pixel-based
DAPS 0.216 (0.045) 27.92 (1.90) 0.154 (0.030) 30.13 (1.70) 0.197 (0.032) 28.64 (1.69)
DiffStateGrad-DAPS (ours) 0.180 (0.031) 29.02 (1.86) 0.119 (0.023) 31.74 (1.57) 0.181 (0.028) 29.35 (1.70)
Latent
PSLD 0.357 (0.060) 22.87 (1.67) 0.322 (0.048) 24.25 (1.47) 0.313 (0.098) 24.51 (2.98)
DiffStateGrad-PSLD (ours) 0.355 (0.043) 22.95 (1.32) 0.319 (0.039) 24.31 (1.18) 0.320 (0.079) 24.56 (2.46)
ReSample 0.253 (0.039) 27.78 (1.79) 0.160 (0.028) 30.55 (1.75) 0.204 (0.038) 28.02 (2.05)
DiffStateGrad-ReSample (ours) 0.245 (0.030) 28.04 (1.66) 0.153 (0.024) 30.82 (1.71) 0.200 (0.031) 28.27 (1.84)
4.3 Performance
We provide quantitative results in Tables 2, 3, 4 and 5, and qualitative results in Figure 9. These results demonstrate the substantial improvement in the performance of DiffStateGrad-PSLD, DiffStateGrad-ReSample, and DiffStateGrad-DAPS against their respective SOTA counterparts across a wide variety of linear and nonlinear tasks on both FFHQ and ImageNet. For example, DiffStateGrad significantly improves performance and increases reconstruction consistency in both pixel-based and latent solvers for phase retrieval (Table 4). We additionally show that DiffStateGrad does not impact the diversity of the posterior sampling (see Figure 15). Finally, we note that DiffStateGrad is robust to the choice of subspace rank (Figures 10 and 8). We refer the reader to Appendix B for the extensive qualitative performance of DiffStateGrad.
Table 4: Performance comparison for nonlinear tasks on FFHQ 256 $\times$ 256.
@@ 4.3 Performance
(a) PSNR
(b) SSIM
(c) LPIPS
Figure 8: Performance of DiffStateGrad-DAPS for different variance retention thresholds $\tau$. DiffStateGrad is robust to the choice of $\tau$, as values $\geq 0.6$ show similar performance in this figure. In our main experiments, we use $\tau=0.99$ (see Section C.1 for further details). We evaluate images from FFHQ $256\times 256$ on box inpainting.
@@ 5 Conclusion
We introduce a Diffusion State-Guided Projected Gradient (DiffStateGrad) to enhance the performance and robustness of diffusion models in solving inverse problems. DiffStateGrad addresses the introduction of artifacts and deviations from the data manifold by constraining gradient updates to a subspace approximating the manifold. DiffStateGrad is versatile, applicable across various diffusion models and sampling algorithms, and includes an adaptive rank that dynamically adjusts to the gradient’s complexity. Overall, DiffStateGrad reduces the need for excessive tuning of hyperparameters and significantly boosts performance for more challenging inverse problems.
We note that DiffStateGrad assumes that the learned prior is a relatively good prior for the task at hand. Since DiffStateGrad encourages the process to stay close to the manifold structure captured by the generative prior, it may introduce the prior’s biases into image restoration tasks. Hence, DiffStateGrad may not be recommended for certain inverse problems such as black hole imaging (Feng et al., 2024).
(a) Box inpainting.
(b) Random inpainting.
(c) Gaussian deblur.
(d) Super-resolution (×$\times 4$).
@@ 6 Reproducibility Statement
To ensure the reproducibility of our results, we thoroughly detail the hyperparameters used in our experiments in Section C.1. We also provide specific implementation and configuration details of all the baselines used in Sections C.3, C.4 and C.5. Moreover, we use easily accessible pre-trained diffusion models throughout our experiments. PSLD (https://github.com/LituRout/PSLD), ReSample (https://github.com/soominkwon/resample), and DAPS (https://github.com/zhangbingliang2019/DAPS) all have publicly available code. We will make the anonymous GitHub repository with our code available to the reviewers after the discussion forum opens.
@@ Appendix A Additional Results
This section provides additional results.
Table 6: Performance comparison for inpainting tasks on ImageNet 256 $\times$ 256.
Method Inpaint (Box) Inpaint (Random)
[…]
@@
Table 8: SSIM comparison for deblurring and super-resolution tasks on FFHQ 256 $\times$ 256.
Method Gaussian deblur Motion deblur SR (x4)
@@
DAPS 0.786 (0.051) 0.837 (0.040) 0.797 (0.044)
DiffStateGrad-DAPS (ours) 0.803 (0.044) 0.853 (0.028) 0.801 (0.039)
@@
PSLD 0.537 (0.094) 0.615 (0.075) 0.650 (0.140)
DiffStateGrad-PSLD (ours) 0.542 (0.077) 0.620 (0.065) 0.640 (0.123)
ReSample 0.757 (0.049) 0.854 (0.034) 0.790 (0.048)
DiffStateGrad-ReSample (ours) 0.767 (0.041) 0.860 (0.031) 0.795 (0.044)
@@ Appendix B Visualizations
This section contains additional visualizations for performance comparison.
@@ C.1 Hyperparameters
For all main experiments across all three methods, we use the variance retention threshold $\tau=0.99$. For all experiments involving PSLD and DAPS, we perform the DiffStateGrad projection step every iteration ($F=1$). For all experiments involving ReSample, we perform the step every five iterations ($F=5$). See the sections dedicated to each method for further implementation details. We reiterate that various values of $\tau$ are reasonable options for optimal performance (see Figures 10 and 8).
@@ C.3 PSLD
Our DiffStateGrad-PSLD algorithm integrates the state-guided projected gradient directly into the PSLD update process. For each iteration (i.e., frequency $F=1$) of the main loop, after computing the standard PSLD update ${\bm{z}}^{\prime}_{t-1}$, we introduce our DiffStateGrad method. First, we calculate the full gradient $\bm{G}_{t}$ according to PSLD, combining both the measurement consistency term and the fixed-point constraint. We then perform SVD on the current latent representation (or diffusion state) ${\bm{Z}}_{t}$ (${\bm{z}}_{t}$ in image matrix form). Using the variance retention threshold $\tau$, we determine the appropriate rank for our projection. We construct projection matrices from the truncated singular vectors and use these to approximate the gradient. This approximated gradient $\bm{G}^{\prime}_{t}$ is then used for the final update step, replacing the separate gradient updates in standard PSLD. This process is repeated at every iteration, allowing for adaptive, low-rank updates throughout the entire diffusion process.

### v2

@@ 4 Results
Experimental setup. We evaluate the performance of DiffStateGrad applied to four SOTA diffusion methods of PSLD (Rout et al., 2023), ReSample (Song et al., 2023a), DPS (Chung et al., 2023), and DAPS (Zhang et al., 2024). These methods span both latent solvers (PSLD and ReSample) and pixel-based solvers (DPS and DAPS). We also directly compare against other methods including DDNM (Wang et al., 2023), DDRM (Kawar et al., 2022), MCG (Chung et al., 2022a), and MPGD-AE (He et al., 2024). We evaluate performance based on key quantitative metrics, including LPIPS (Learned Perceptual Image Patch Similarity), PSNR (Peak Signal-to-Noise Ratio), and SSIM (Structural Similarity Index) (Wang et al., 2004). We demonstrate the effectiveness of DiffStateGrad on two datasets: a) the FFHQ $256\times 256$ validation dataset (Karras et al., 2021), and b) the ImageNet $256\times 256$ validation dataset (Deng et al., 2009). For pixel-based experiments, we use (i) the pre-trained diffusion model from (Chung et al., 2023) for the FFHQ dataset, and (ii) the pre-trained model from (Dhariwal & Nichol, 2021) for the ImageNet dataset. For latent diffusion experiments, we use (i) the unconditional LDM-VQ-4 model trained on FFHQ (Rombach et al., 2022) for the FFHQ dataset, and (ii) the Stable Diffusion v1.5 (Rombach et al., 2022) model for the ImageNet dataset.
@@
(b) Motion Deblur.
Figure 7: Robustness of DiffStateGrad to measurement noise. We evaluate the performance of DiffStateGrad-PSLD and DiffStateGrad-DAPS with their respective counterparts on different tasks across a range of measurement noise levels (std of 0 to 0.2) using FFHQ $256\times 256$.
Table 3: Performance comparison for linear and nonlinear tasks on FFHQ 256 $\times$ 256.
Method Inpaint (Box) Inpaint (Random) Gaussian deblur Motion deblur SR (x4)
@@
DAPS 0.136 24.57 0.130 30.79 0.216 27.92 0.154 30.13 0.197 28.64
DiffStateGrad-DAPS (ours) 0.113 24.78 0.099 32.04 0.180 29.02 0.119 31.74 0.181 29.35
DPS 0.127 23.91 0.130 28.67 0.145 25.48 0.132 26.75 0.191 24.38
DiffStateGrad-DPS (ours) 0.114 24.10 0.107 30.15 0.128 26.29 0.118 27.61 0.186 24.65
DDNM 0.235 24.47 0.121 29.91 0.216 28.20 - - 0.197 28.03
DDRM 0.159 22.37 0.218 25.75 0.236 23.36 - - 0.210 27.65
MCG 0.309 19.97 0.286 21.57 0.340 6.72 0.702 6.72 0.520 20.05
MPGD-AE 0.138 21.59 0.172 25.22 0.150 24.42 0.120 25.72 0.168 24.01
@@
PSLD 0.158 24.22 0.246 29.05 0.357 22.87 0.322 24.25 0.313 24.51
ReSample 0.198 19.91 0.115 31.27 0.253 27.78 0.160 30.55 0.204 28.02
DiffStateGrad-ReSample (ours) 0.156 23.59 0.106 31.91 0.245 28.04 0.153 30.82 0.200 28.27
(a) Linear inverse problems.
@@
(b) Nonlinear inverse problems.
(a) Box inpainting.
(b) Random inpainting.
We consider both linear and nonlinear inverse problems for natural images. For evaluation, we sample a fixed set of $100$ images from the FFHQ and ImageNet validation sets. Images are normalized to the range $[0,1]$. We use the default settings for all experiments (see Appendix C for more details). For linear inverse problems, we consider (1) box inpainting, (2) random inpainting, (3) Gaussian deblur, (4) motion deblur, and (5) super-resolution. In the box inpainting task, a random $128\times 128$ box is used, while the random inpainting task employs a $70\%$ random mask. Gaussian and motion deblurring tasks utilize kernels of size $61\times 61$, with standard deviations of $3.0$ and $0.5$, respectively. For super-resolution, images are downscaled by a factor of $4$ using a bicubic resizer. For nonlinear inverse problems, we consider (1) phase retrieval, (2) nonlinear deblur, and (3) high dynamic range (HDR). For phase retrieval, we use an oversampling rate of 2.0, and due to the instability and non-uniqueness of reconstruction, we adopt the strategy from DPS (Chung et al., 2023) and DAPS (Zhang et al., 2024), generating four separate reconstructions and reporting the best result. Like DAPS (Zhang et al., 2024), we normalize the data to lie in the range [0, 1] before applying the discrete Fourier transform. For nonlinear deblur, we use the default setting from (Tran et al., 2021). For HDR, we use a scale factor of 2. We note that PSLD is not designed to handle nonlinear inverse problems. We also conduct an additional experiment for Magnetic Resonance Imaging (MRI) (see Appendix E).
Robustness.Figure 6 exhibits the sensitivity of PSLD to the choice of MG step size; the performance of PSLD significantly deteriorates when a relatively large MG step size is used, leading to poor results across all tasks. In contrast, DiffStateGrad-PSLD shows superior robustness, maintaining high performance over a wide range of MG step sizes. Figure 7 demonstrates the robustness of DiffStateGrad methods compared to their non-DiffStateGrad counterparts when faced with increasing measurement noise. For both inpainting and Gaussian deblur tasks, the performance of DAPS and PSLD deteriorates significantly as noise levels rise. In contrast, DiffStateGrad-DAPS and DiffStateGrad-PSLD exhibit superior resilience across the range of noise levels tested.
Performance. We provide quantitative results in Tables 3(a), 3(b) and 4(b), and qualitative results in Figure 8. These results demonstrate the substantial improvement in the performance of DiffStateGrad-PSLD, DiffStateGrad-ReSample, DiffStateGrad-DPS, and DiffStateGrad-DAPS against their respective SOTA counterparts across a wide variety of linear and nonlinear tasks on both FFHQ and ImageNet. For example, DiffStateGrad significantly improves performance and increases reconstruction consistency in both pixel-based and latent solvers for phase retrieval (Table 3(b)). Our results also highlight the superiority of our choice of subspace compared to prior works. For example, DiffStateGrad-DPS is superior to MCG, which is the special case of DPS for manifold constrained diffusion. Furthermore, the outperformance of DiffStateGrad-DAPS against DAPS and MPGD-AE emphasizes the effectiveness of the proposed SVD-based subspace projection. We additionally show that DiffStateGrad does not impact the diversity of the posterior sampling (see Figure 17). Finally, we note that DiffStateGrad is robust to the choice of subspace rank (Figures 9 and 10). We refer the reader to Appendix B for the extensive qualitative performance of DiffStateGrad.
@@
Method Inpaint (Box) Inpaint (Random)
[…]
Pixel-based
DAPS 0.217 (0.043) 0.762 (0.041) 20.90 (3.69) 0.158 (0.039) 0.794 (0.067) 28.34 (3.65)
DiffStateGrad-DAPS (ours) 0.191 (0.044) 0.801 (0.056) 21.07 (3.77) 0.107 (0.037) 0.856 (0.067) 29.78 (4.17)
DPS 0.257 (0.086) 0.718 (0.097) 19.85 (3.54) 0.256 (0.133) 0.728 (0.143) 26.25 (4.15)
DiffStateGrad-DPS (ours) 0.243 (0.093) 0.731 (0.100) 19.87 (3.61) 0.233 (0.138) 0.754 (0.150) 27.28 (4.88)
MPGD-AE 0.295 (0.057) 0.621 (0.053) 16.12 (2.26) 0.554 (0.148) 0.388 (0.112) 17.91 (3.25)
Latent
(a) Inpainting inverse problems.
@@
Pixel-based
DAPS 0.266 (0.087) 25.27 (3.56) 0.166 (0.058) 28.85 (3.64) 0.259 (0.073) 25.67 (3.40)
DiffStateGrad-DAPS (ours) 0.243 (0.075) 25.87 (3.56) 0.143 (0.050) 29.71 (3.54) 0.229 (0.057) 26.40 (3.44)
MCG 0.550 (-) 16.32 (-) 0.758 (-) 5.89 (-) 0.637 (-) 13.39 (-)
Latent
@@
(b) Deblurring and super-resolution inverse problems.
@@ 5 Conclusion
We introduce a Diffusion State-Guided Projected Gradient (DiffStateGrad) to enhance the performance and robustness of diffusion models in solving inverse problems. DiffStateGrad addresses the introduction of artifacts and deviations from the data manifold by constraining gradient updates to a subspace approximating the manifold. DiffStateGrad is versatile, applicable across various diffusion models and sampling algorithms, and includes an adaptive rank that dynamically adjusts to the gradient’s complexity. Overall, DiffStateGrad reduces the need for excessive tuning of hyperparameters and significantly boosts performance for more challenging inverse problems. We note that DiffStateGrad assumes that the learned prior is a relatively good prior for the task at hand. Since DiffStateGrad encourages the process to stay close to the manifold structure captured by the generative prior, it may introduce the prior’s biases into image restoration tasks. Hence, DiffStateGrad may not be recommended for certain inverse problems such as black hole imaging (Feng et al., 2024). We finally note that DiffStateGrad can be combined with prior works that adopt initialization strategies for the diffusion process to further accelerate and improve performance (Fabian et al., 2024; Chung et al., 2022b). We leave this for future work.
@@ 6 Reproducibility Statement
To ensure the reproducibility of our results, we thoroughly detail the hyperparameters used in our experiments in C.1. We also provide specific implementation and configuration details of all the baselines used in C.3C.4C.5, and C.6. Moreover, we use easily accessible pre-trained diffusion models throughout our experiments. PSLD (https://github.com/LituRout/PSLD), ReSample (https://github.com/soominkwon/resample), DPS (https://github.com/DPS2022/diffusion-posterior-sampling), and DAPS (https://github.com/zhangbingliang2019/DAPS) all have publicly available code. We will make our code available upon acceptance to the conference.
@@ Appendix A Additional Results
(a) PSNR
(b) SSIM
(c) LPIPS
Figure 10: Performance of DiffStateGrad-DAPS for different variance retention thresholds $\tau$. DiffStateGrad is robust to the choice of $\tau$, as values $\geq 0.6$ show similar performance in this figure. In our main experiments, we use $\tau=0.99$ (see C.1 for further details). We evaluate images from FFHQ $256\times 256$ on box inpainting.
(a) PSNR
(b) SSIM
(c) LPIPS
Figure 11: Performance of DiffStateGrad-DAPS for different projection periods $P$. As period of projection increases, DiffStateGrad still outperforms DAPS without projection, which is the SOTA. See C.1 for details of $P$ values used in our experiments. We evaluate images from FFHQ $256\times 256$ on random inpainting.
Table 6: SSIM comparison for linear tasks on FFHQ 256 $\times$ 256.
Method Inpaint (Box) Inpaint (Random) Gaussian deblur Motion deblur SR (x4)
@@
DAPS 0.806 (0.028) 0.829 (0.022) 0.786 (0.051) 0.837 (0.040) 0.797 (0.044)
DiffStateGrad-DAPS (ours) 0.849 (0.029) 0.887 (0.023) 0.803 (0.044) 0.853 (0.028) 0.801 (0.039)
DPS 0.810 (0.039) 0.815 (0.045) 0.709 (0.062) 0.754 (0.056) 0.675 (0.071)
DiffStateGrad-DPS (ours) 0.831 (0.039) 0.852 (0.046) 0.739 (0.062) 0.782 (0.056) 0.683 (0.073)
MPGD-AE 0.753 (0.029) 0.731 (0.050) 0.664 (0.071) 0.723 (0.061) 0.670 (0.070)
@@
PSLD 0.819 (0.031) 0.809 (0.049) 0.537 (0.094) 0.615 (0.075) 0.650 (0.140)
DiffStateGrad-PSLD (ours) 0.880 (0.028) 0.898 (0.024) 0.542 (0.077) 0.620 (0.065) 0.640 (0.123)
ReSample 0.807 (0.036) 0.892 (0.030) 0.757 (0.049) 0.854 (0.034) 0.790 (0.048)
DiffStateGrad-ReSample (ours) 0.841 (0.032) 0.913 (0.023) 0.767 (0.041) 0.860 (0.031) 0.795 (0.044)
@@ C.1 Hyperparameters
For all main experiments across all four methods, we use the variance retention threshold $\tau=0.99$. For all experiments involving PSLD, DPS, and DAPS, we perform the DiffStateGrad projection step every iteration ($P=1$). For all experiments involving ReSample, we perform the step every five iterations ($P=5$). See the sections dedicated to each method for further implementation details. We reiterate that various values of $\tau$ and $P$ are reasonable options for optimal performance (see Figures 9, 10 and 11).
@@ C.3 PSLD
Our DiffStateGrad-PSLD algorithm integrates the state-guided projected gradient directly into the PSLD update process. For each iteration (i.e., period $P=1$) of the main loop, after computing the standard PSLD update ${\bm{z}}^{\prime}_{t-1}$, we introduce our DiffStateGrad method. First, we calculate the full gradient $\bm{G}_{t}$ according to PSLD, combining both the measurement consistency term and the fixed-point constraint. We then perform SVD on the current latent representation (or diffusion state) ${\bm{Z}}_{t}$ (${\bm{z}}_{t}$ in image matrix form). Using the variance retention threshold $\tau$, we determine the appropriate rank for our projection. We construct projection matrices from the truncated singular vectors and use these to approximate the gradient. This approximated gradient $\bm{G}^{\prime}_{t}$ is then used for the final update step, replacing the separate gradient updates in standard PSLD. This process is repeated at every iteration, allowing for adaptive, low-rank updates throughout the entire diffusion process.